# From Zipf‘s Law to Neural Scaling through Heaps‘ Law and Hilberg‘s Hypothesis

### v1

@@ 3.1 Zipf‘s law implies Heaps‘ law
Let us recall from Section 2.4 that the expected number of hapaxes for IID processes satisfies
$V(t+1|1)=(t+1)\Delta V(t).$ (64)
The growth rate of the number of hapaxes $V(t+1|1)$ is controlled by the tail of the marginal distribution. In fact, there are two simple symmetric conditions on the tail of the marginal distribution that sandwich bound the number of hapaxes and lead to a differential Heaps law.
We continue notation $p_{k}:=P(K_{t}=k)$ in the following.
@@
For an IID process $(K_{t})_{t\in\mathbb{Z}}$, suppose that
$\sum_{k}{\bf 1}{\left\{p_{k}\geq p\right\}}\leq C_{1}p^{-\beta}$ (65)
for some $\beta\in(0,1)$, $C_{1}>0$, and all $p>0$. Then
$\Delta V(t)\leq C_{2}t^{\beta-1},$ (66)
where $C_{2}:=\Gamma(1-\beta)C_{1}$.
@@
We may bound
$\Delta V(t)$ $=\frac{V(t+1|1)}{t+1}=\sum_{k}p_{k}(1-p_{k})^{t}$
$\leq\sum_{k}\int_{0}^{p_{k}}(1-p)^{t}dp=\int_{0}^{1}\left(\sum_{k}{\bf 1}{\left\{p_{k}\geq p\right\}}\right)(1-p)^{t}dp$
$\leq C_{1}\int_{0}^{1}p^{-\beta}(1-p)^{t}dp.$ (67)
@@
The reverse bound can be demonstrated for a larger class of processes. For this aim, let us define a suitable mixing coefficient.
Definition 5.
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$, denote the mixing coefficient
$\Psi:=\sup_{t,k}\frac{\sum_{i=1}^{t}\sum_{j=1}^{t}P(K_{i}=k,K_{j}=k)-tp_{k}}{t(t-1)p_{k}^{2}}.$ (69)
In particular, for an IID process, we have $\Psi=1$. For a stationary ergodic process, we have $\lim_{t\to\infty}t^{-1}\sum_{i=1}^{t}P(K_{0}=k,K_{i}=k)=p_{k}^{2}$ [26, Theorem 4.9]. For a comprehensive survey of various mixing coefficients and mixing conditions, we refer to [12]. For our purposes, it is sufficient to assume that $\Psi$ is finite.
@@
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ with $\Psi<\infty$, suppose that
$\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}\geq C_{3}p^{1-\beta}$ (70)
holds for some $\beta\in(0,1)$, $C_{3}>0$, and all $p>0$. Then we have
@@
$=\sum_{k}\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)$
$=\sum_{k}\max\left\{0,\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)\right\}$
$\geq\sum_{k}\max\left\{0,2p_{k}-\frac{1}{2t}\sum_{i=1}^{2t}\sum_{j=1}^{2t}P(K_{i}=k,K_{j}=k)\right\}$
$\geq\sum_{k}p_{k}\max\left\{0,1-\Psi(2t-1)p_{k}\right\}\geq\sum_{k}p_{k}\max\left\{0,1-2\Psi tp_{k}\right\}$
$\geq\frac{1}{2}\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\leq\frac{1}{4\Psi t}\right\}}\geq\frac{C_{3}(4\Psi t)^{\beta-1}}{2}.$ (72)
Now we can show that conditions (65) and (70) hold under a broadly understood Zipf law.
@@
$C_{5}k^{-\alpha}\leq p_{k}\leq C_{6}k^{-\alpha}$ (73)
for some $\alpha>1$, $C_{5},C_{6}>0$, and all $k\in\mathbb{N}$. Then for $\beta=1/\alpha$, we have
$\sum_{k}{\bf 1}{\left\{p_{k}\geq p\right\}}$ $\leq\int_{0}^{\infty}{\bf 1}{\left\{C_{6}k^{-\alpha}\geq p\right\}}dk$
@@
$\sum_{k}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}$ $\geq\int_{1}^{\infty}C_{5}k^{-\alpha}{\bf 1}{\left\{C_{6}k^{-\alpha}\leq p\right\}}dk$
$=\int_{1}^{\infty}C_{5}k^{-\alpha}{\bf 1}{\left\{k\geq\left(\frac{p}{C_{6}}\right)^{-\beta}\right\}}dk=\frac{C_{5}}{1/\beta-1}\left(\frac{p}{C_{6}}\right)^{1-\beta}.$ (75)
Thus conditions (65) and (70) hold for some constants $C_{1},C_{3}>0$. In consequence, an approximate Zipf law (73) implies the differential Heaps law (66) if the process is IID and it implies the differential Heaps law (71) if the process is stationary and $\Psi<\infty$.
@@ 3.2 Heaps‘ law implies Hilberg‘s law
Now we will formalize the remark made in Section 2.4. Namely, we will use the correspondence that the expected cardinality applies to arbitrary (stationary) processes as the Shannon entropy applies to (stationary) Santa Fe processes. In particular, the differential Heaps law (71) for a stationary narration implies a differential Hilberg law.
@@
holds in particular if the alphabet of process $(K_{t})_{t\in\mathbb{Z}}$ is finite (see [26, Sections 5.3 and 5.4] on a generalization of Shannon information measures to arbitrary $\sigma$-fields) or if $H(K_{t})<\infty$ (this follows by approximating $K_{t}$ by finite-alphabet random variables). Now suppose that process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and $H(K_{t})<\infty$. We have $h=H(K_{t+1}|K_{-\infty}^{t})$ by (82) and $v=0$ by Proposition 9. Since $H(K_{k+1}^{k+s}|K_{1}^{t})\geq H(K_{k+1}^{k+s}|K_{-\infty}^{k})=hs$ for $k\geq t$, we may write
@@ 4 Conclusion
Formalizing prior results in machine learning [50, 61, 64, 14, 70, 71] and extending earlier results in information theory and quantitative linguistics [23, 25, 26, 28], we have derived of a chain of deductive implications that connect Zipf‘s law to neural scaling. By isolating the discrete steps that go through Heaps‘ law and Hilberg‘s hypothesis and by giving explicit assumptions needed for each step, we have attempted to clarify which aspects of natural language properties are responsible for the power-law behavior observed in foundation models. Our results show that once vocabulary growth exhibits a power-law growth and once block entropy inherits this scaling, then the constraints imposed by limited data, parameters, and compute naturally produce a generic neural scaling law.
Our theoretical derivation of underparameterization as the optimal regime contrasts sharply with the empirical success of overparameterized models [51]. This discrepancy suggests that naive parameter counts may overestimate the effective capacity of real-valued neural networks, as larger models may use fewer digits of the binary expansion per real-valued parameter. If this holds true then the received concept of overparameterization may be misleading and determining the true number of degrees of freedom of a foundation model may not be so straightforward. Likewise, our purely information-theoretic modeling of compute—via entropy constraints—offers only a coarse abstraction of training dynamics. We suppose that a refined treatment may require tools from resource-bounded algorithmic information theory.
Finally, the reduction of the neural scaling law to Hilberg‘s hypothesis shifts attention to the deeper question of why power-law scaling of entropy arises in natural language at all. One simple possibility is that Hilberg‘s law emerges for Zipfian vocabularies with word shapes that are sufficiently arbitrary or algorithmically random, as suggested by the Santa Fe decomposition (7). Establishing such a connection rigorously requires a more detailed analysis of universal coding, extending those of works [25, 26, 27]. We hope that the framework developed here provides a baseline for such investigations and that future work will refine the idealizations that we have identified.
AI assistance
Several paragraphs in the introduction and in the conclusion have been co-drafted by the author and ChatGPT and critically post-edited by the author. No reference was suggested by the AI assistant.
Funding

### v2

@@ 3.1 Zipf‘s law implies Heaps‘ law
By formulas (58) and (63), the expected rate of hapaxes for IID processes equals
$\frac{V(t|1)}{t}=\Delta V(t-1),$ (65)
which can be contrasted with the more general inequality (54) for stationary sources. Thus some bounds for difference $\Delta V(t)$ can be obtained by showing that the rate of hapaxes $V(t|1)/t$ is controlled by the tail of the marginal distribution. In fact, there are two symmetric conditions on the tail of the marginal distribution that sandwich the number of hapaxes and lead to a differential Heaps law. Similar bounds, though asymptotic and covering only the IID case, were discussed by Karlin [54].
In the following, we use $p_{k}:=P(K_{0}=k)$ and $p_{k}(t):=P(K_{t}=k|K_{0}=k)$.
@@
For an IID process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers, suppose that
$\sum_{k\in\mathbb{N}}{\bf 1}{\left\{p_{k}\geq p\right\}}\leq C_{0}p^{-\beta}$ (66)
for some $\beta\in(0,1)$, $C_{0}>0$, and all $p>0$. Then
$\Delta V(t)\leq C_{1}t^{\beta-1},$ (67)
where $C_{1}:=\Gamma(1-\beta)C_{0}$.
@@
By (65) and (62), we may bound
$\Delta V(t)$ $=\frac{V(t+1|1)}{t+1}=\sum_{k\in\mathbb{N}}p_{k}(1-p_{k})^{t}$
$\leq\sum_{k\in\mathbb{N}}\int_{0}^{p_{k}}(1-p)^{t}dp=\int_{0}^{1}\left(\sum_{k\in\mathbb{N}}{\bf 1}{\left\{p_{k}\geq p\right\}}\right)(1-p)^{t}dp$
$\leq C_{0}\int_{0}^{1}p^{-\beta}(1-p)^{t}dp.$ (68)
@@
The reverse bound can be demonstrated for a larger class of processes.
@@
For a stationary process $(K_{t})_{t\in\mathbb{Z}}$ over natural numbers, suppose that
$\sum_{k\in\mathbb{N}}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}$ $\geq C_{2}p^{1-\beta},$ (70)
$\frac{p_{k}(t)}{p_{k}}$ $\leq C_{3}.$ (71)
holds for some $\beta\in(0,1)$, $C_{2},C_{3}>0$, all $k,t\in\mathbb{N}$, and all $p>0$. Then we have
@@
$=\sum_{k\in\mathbb{N}}\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)$
$=\sum_{k\in\mathbb{N}}\max\left\{0,\frac{1}{2t}\sum_{i=1}^{2t}P\left(K_{i}=k\not\in\mathcal{K}_{1}^{i-1}\cup\mathcal{K}_{i+1}^{2t}\right)\right\}$
$\geq\sum_{k\in\mathbb{N}}\max\left\{0,2p_{k}-\frac{1}{2t}\sum_{i=1}^{2t}\sum_{j=1}^{2t}P(K_{i}=k,K_{j}=k)\right\}$
$\geq\sum_{k\in\mathbb{N}}p_{k}\max\left\{0,1-C_{3}(2t-1)p_{k}\right\}\geq\sum_{k\in\mathbb{N}}p_{k}\max\left\{0,1-2C_{3}tp_{k}\right\}$
$\geq\frac{1}{2}\sum_{k\in\mathbb{N}}p_{k}{\bf 1}{\left\{p_{k}\leq\frac{1}{4C_{3}t}\right\}}\geq\frac{C_{2}(4C_{3}t)^{\beta-1}}{2}.$ (73)
We can show that conditions (66) and (70) hold under a broadly understood Zipf law.
@@
$C_{5}k^{-1/\beta}\leq p_{k}\leq C_{6}k^{-1/\beta}$ (74)
for some $\beta\in(0,1)$, $C_{5},C_{6}>0$, and all $k\in\mathbb{N}$. Then we have
$\sum_{k\in\mathbb{N}}{\bf 1}{\left\{p_{k}\geq p\right\}}$ $\leq\int_{0}^{\infty}{\bf 1}{\left\{C_{6}k^{-1/\beta}\geq p\right\}}dk$
@@
$\sum_{k\in\mathbb{N}}p_{k}{\bf 1}{\left\{p_{k}\leq p\right\}}$ $\geq\int_{1}^{\infty}C_{5}k^{-1/\beta}{\bf 1}{\left\{C_{6}k^{-1/\beta}\leq p\right\}}dk$
$=\int_{1}^{\infty}C_{5}k^{-1/\beta}{\bf 1}{\left\{k\geq\left(\frac{p}{C_{6}}\right)^{-\beta}\right\}}dk=\frac{C_{5}}{1/\beta-1}\left(\frac{p}{C_{6}}\right)^{1-\beta}.$ (76)
Thus conditions (66) and (70) hold for some constants $C_{0},C_{2}>0$. In consequence, an approximate Zipf law (74) implies the differential Heaps law (67) if the process is IID and, combined with (71), it implies the differential Heaps law (72) if the process is stationary.
We suppose that the above results can be somewhat generalized by applying or developing classical results by Karlin [54].
@@ 3.2 Heaps‘ law implies Hilberg‘s law
Now we will formalize the remark made in Section 2.4. Namely, we will use the correspondence that the expected cardinality applies to arbitrary (stationary) processes analogously as the Shannon entropy applies to (stationary) Santa Fe processes. In particular, the differential Heaps law (72) for a stationary narration implies a differential Hilberg law.
@@
holds in particular if the alphabet of process $(K_{t})_{t\in\mathbb{Z}}$ is finite (see [28, Sections 5.3 and 5.4] on a generalization of Shannon information measures to arbitrary $\sigma$-fields) or if $H(K_{t})<\infty$ (this follows by approximating $K_{t}$ by finite-alphabet random variables). Now suppose that process $(K_{t})_{t\in\mathbb{Z}}$ is stationary and $H(K_{t})<\infty$. We have $h=H(K_{t+1}|K_{-\infty}^{t})$ by (83) and $v=0$ by Proposition 9. Since $H(K_{k+1}^{k+s}|K_{1}^{t})\geq H(K_{k+1}^{k+s}|K_{-\infty}^{k})=hs$ for $k\geq t$, we may write
@@ 4 Conclusion
Formalizing prior results in machine learning [52, 65, 68, 15, 74, 75] and extending earlier results in information theory and quantitative linguistics [25, 27, 28, 30], we have derived of a deductive chain that connects Zipf‘s law to neural scaling. By isolating the discrete steps that go through Heaps‘ law and Hilberg‘s hypothesis and by giving explicit assumptions needed for each step, we have attempted to clarify which aspects of natural language are responsible for the power-law behavior observed in foundation models. Our results show that once vocabulary growth exhibits a power-law growth and once block entropy inherits this scaling, then the constraints imposed by limited data, parameters, and compute produce the neural scaling law.
Our theoretical derivation of underparameterization as the optimal regime contrasts sharply with the empirical success of overparameterized models [53, 50, 76]. This discrepancy suggests that naive parameter counts may overestimate the effective capacity of real-valued neural networks, as larger models may use fewer digits of the binary expansion per real-valued parameter. If this holds true then the received concept of overparameterization may be misleading and determining the true number of degrees of freedom of a foundation model may not be so straightforward. Likewise, our purely information-theoretic modeling of compute—via entropy constraints—offers only a coarse abstraction of training dynamics. We suppose that a refined treatment may rest on resource-bounded algorithmic information theory [59, 60].
Finally, the reduction of the neural scaling law to Hilberg‘s hypothesis shifts attention to the deeper question of why power-law scaling of entropy arises in natural language at all. One simple possibility is that Hilberg‘s law is equvalent to Zipf‘s law if word shapes are sufficiently arbitrary or algorithmically random, as suggested by the Santa Fe decomposition (7). Establishing such a connection rigorously requires a more detailed analysis of universal coding, extending those of works [27, 28, 29]. We hope that the framework developed here provides a baseline for such investigations and that future work will refine the idealizations that we have identified.
Acknowledgments and Disclosure of Funding
Several paragraphs in the Introduction and in the Conclusion were drafted in a dialog between the author and ChatGPT (https://chatgpt.com/) and critically post-edited by the author. Subsequently, the article was reviewed by the Stanford Agentic Reviewer (https://paperreview.ai/) and its suggestions were partially followed by the author.